OpenStudy (anonymous):

Solve for x: 2 over 5 (x - 2) = 4x. 2 over 9 9 negative 2 over 9 negative 9 over 2

OpenStudy (anonymous):

what?????

OpenStudy (anonymous):

i think this could be the answer.. 2/5 ( x - 2 ) = 4x 0.4 ( x - 2 ) = 4x 0.4x - 0.8 = 4x -0.8 = 4x - 0.4 >> -0.4+4 -0.8 = 3.6x -0.8 / 3.6 = 3.6x / 3.6 -0.22 = x >> OR >> x = - 2/9 This is my answer..

OpenStudy (texaschic101):

2/5(x - 2) = 4x --- distribute through the parenthesis 2/5x - 4/5 = 4x --- multiply by 5 2x - 4 = 20x -4 = 20x - 2x -4 = 18x -4/18 = x -2/9 = x I agree with Maxlure :)

OpenStudy (anonymous):

@texaschic101 hhahhahah thanks..

OpenStudy (texaschic101):

:)